An airplane starts from rest and accelerates on a runway at 5m/s^2 for 7 seconds. How far down the runway is the plane after 7 seconds?
1
Expert's answer
2019-10-22T10:21:27-0400

The distance traveled by airplane


"d=v_it+\frac{at^2}{2}=0\times 7+\frac{5\times 7^2}{2}=122.5\;\rm m"